# Standard Gaussian Process is All You Need for High-Dimensional Bayesian Optimization

### v2

@@ 4.2.3 Computational Efficiency
Figure 4: Alignment between the prediction (y-axis) and ground-truth function values (x-axis). The top row is the results for Mopta08 (124) and the second row DNA (180).
@@ 5 Conclusion
A longstanding belief suggests that standard Gaussian process is insufficient to model high-dimensional functions, and Bayesian optimization with standard GP is expected to exhibit poor performance when optimizing functions with more than 20 variables. However, through a comprehensive investigation across various benchmark tasks, this paper challenges and overturns the prevailing notion.
Our findings reveal that standard GP excels in function estimation and uncertainty calibration, even when dealing with hundreds of input variables. Bayesian Optimization (BO) conducted with standard GP demonstrates remarkable performance in high-dimensional optimization, frequently outperforming existing high-dimensional BO methods by a substantial margin. Those methods, incorporating strong structural assumptions and employing more complex surrogate modeling, exhibit less robustness to various target functions, whose intrinsic structural properties are often agnostic in practice. Furthermore, the standard GP, benefiting from a parsimonious parameter space, can achieve satisfactory performance using only maximum likelihood estimation, thus eliminating the need for more computationally expensive Monte-Carlo methods in practical applications. Given these revelations, we advocate for a comprehensive exploration of the potential of the standard GP in high-dimensional optimization tasks.
Appendix
A More Experiment Details
@@ A.1 Definition of Synthetic Functions
Ackley. We used the Ackley function defined in (Surjanovic and Bingham, ),
@@
where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in (Surjanovic and Bingham, ). The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
@@
Figure 5: Maximum Function Value Obtained vs. Number of Steps in synthetic benchmarks. The results were averaged from 10 runs.
@@
Figure 6: Maximum Function Value Obtained vs. Number of Steps in real-world benchmarks. The results were averaged from 10 runs.
A.2 NN-Prune Details
Given specific per-layer pruning rates ${\bf x}=[x_{1},\ldots,x_{151}]$, the pruning procedure is as follows. The weights in each layer $i$ are first ranked according to their magnitudes, and the bottom $x_{i}$ portion of the weights are removed. The pruned network is then fine tuned on Tiny ImageNet (Le and Yang, 2015). The prediction accuracy after pruning and fine tuning on a separate test dataset is used as the target function. To respect the global pruning rate constraint $\alpha$, we add to the objective a soft penalty term $-\lambda\cdot\max(\alpha-\beta,0)$, where $\beta$ is the global pruning rate reflected from ${\bf x}$. Note that the penalty is only incurred when $\alpha>\beta$. When $\alpha<\beta$, namely, the network is pruned to be sparser than the expected level, the test accuracy typically experienced a significant decline. Consequently, the optimization process automatically adjusts to increase $\beta$. We set $\lambda=8.0$ to introduce a strong penalty. In our experiment, we employed $\alpha=0.5$.
A.3 Implementation
A.4 Computing Environment
@@
We conducted the experimental investigations on a large computer cluster equipped with Intel Cascade Lake Platinum 8268 chips. For the NN-Prun problem, we ran the experiments on a Linux workstation, with a 16-core Intel 13900K CPU and one NVIDIA RTX 4090 GPU.
Task GP-ARD GP SaasBO ALEBO-10 ALEBO-20
Mopta08(124) 0.048$\pm$0.015 0.126$\pm$0.041 0.052$\pm$0.018 0.654$\pm$0.146 0.659$\pm$0.096
Rover(60) 0.493$\pm$0.126 0.449$\pm$0.108 0.453$\pm$0.121 0.813$\pm$0.168 1.266$\pm$0.352
DNA(180) 0.098$\pm$0.028 0.175$\pm$0.070 0.115$\pm$0.039 0.827$\pm$0.173 0.955$\pm$0.335
SVM(388) 0.042$\pm$0.010 0.098$\pm$0.012 0.061$\pm$0.017 0.624$\pm$0.159 0.493$\pm$0.132
Ackley(150, 150) 0.010$\pm$0.003 0.004$\pm$0.001 0.004$\pm$0.001 0.367$\pm$0.074 0.102$\pm$0.015
Ackley(300, 150) 0.010$\pm$0.003 0.018$\pm$0.004 0.010$\pm$0.002 0.360$\pm$0.109 0.177$\pm$0.075
Hartmann6(300, 6) 0.020$\pm$0.009 0.769$\pm$0.124 0.077$\pm$0.045 1.012$\pm$0.154 1.286$\pm$0.210
Rosenbrock(100, 100) 0.049$\pm$0.006 0.124$\pm$0.020 0.062$\pm$0.008 0.601$\pm$0.095 0.807$\pm$0.248
Rosenbrock(300, 100) 0.048$\pm$0.012 0.202$\pm$0.035 0.073$\pm$0.021 0.568$\pm$0.057 1.290$\pm$0.882
Stybtang(200, 200) 0.040$\pm$0.014 0.078$\pm$0.012 0.044$\pm$0.014 0.502$\pm$0.066 0.861$\pm$0.329
@@
Table 5: Mean Squared Error (MSE) of each surrogate model. The results were averaged over five runs.
@@
Figure 7: Optimization performance when using GP-ARD trained with maximum likelihood estimation (MLE) and NUTS on the rest of benchmarks. The results were summarized from 10 runs.

### v3

@@ 4.3 Surrogate Learning
Figure 3: Alignment between the prediction (y-axis) and ground-truth function values (x-axis). The top row includes the results for Mopta08 (124) and the second row DNA (180).
We repeated the experiment for five times, and evaluated mean-square-error (MSE), the average test log-likelihood (LL), and the standard deviation. Table 4 and Appendix Table 5 report the results for all the GP models. The best and second best results are highlighted with a bold font. As we can see, in every dataset, standard GP with Mateŕ kernels achieves either the best or the second best result, demonstrating excellent and robust performance in learning high-dimensional functions.
In addition, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180). As illustrated in Fig. 3, the prediction (and the predictive standard deviation) of ARD Matérn matches the ground-truth equally as well as SaasBO-NUTS, with a slight improvement over Matérn (using a single length-scale parameter).
Together these results demonstrate that directly incorporating high-dimensional inputs into the kernel function does not necessarily compromise the learning performance. This implies that, even when the input space is high dimensional, the popular Matérn kernel remains expressive enough to capture the similarities in the space, allowing them to perform well in covariance modeling. On the contrary, enforcing (unnecessary) low-rank structures or sparsity might cause information loss, potentially detrimental to optimization performance. Our observation is consistent with the recent work in kernel operator learning [3], where the standard kernel (GP) regression has demonstrated the state-of-the-art performance in learning function-to-function mappings. In this context, the input comprises of a discretized function, typically of very high-dimensionality (e.g., several thousands).
4.4 Computational Efficiency
So far, in our BO loop, the GP model is trained via MAP or MLE. Although the performance has been promising, it is worth investigating whether a full Bayesian approach would yield substantial improvements. In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction. We conducted another round of BO experiments. For each step, we used NUTS to train GP with ARD Matérn. We adopted the same settings as used in SaasBO-NUTS [6]: The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples. We show the running and final optimization performance in Fig 4 , and Appendix Fig. 9. It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal. However, running NUTS is much more computationally costly compared to doing MAP and MLE. For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step. Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient. It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
By contrast, for more complex surrogate models, such as SaasBO, the MCMC based training might be more crucial for achieving good performance. As depicted in Fig. 2, SaasBO-NUTS demonstrates a substantial improvement over SaasBO-MAP across all real-world benchmarks, excluding NAS201. However, the computational cost of running SaasBO-NUTS is considerably higher, ranging from 5.4x (NAS201) to 33.9x (SVM, DNA) compared to the running time of SaasBO-MAP. The substantial increase in computational costs may pose a bottleneck for practical usage. Detailed per-optimization-step running time for these methods are provided in Appendix Table 6 and 7.
Dataset ARD Matérn Matérn SaasBO Proj-GP Add-GP
Mopta08(124) 0.034$\pm$0.006 0.145$\pm$0.057 0.038$\pm$0.008 0.461$\pm$0.157 0.107$\pm$0.033
Rover(60) 0.327$\pm$0.078 0.284$\pm$0.09 0.284$\pm$0.09 0.548$\pm$0.097 0.321$\pm$0.097
DNA(180) 0.149$\pm$0.038 0.332$\pm$0.053 0.152$\pm$0.049 0.747$\pm$0.121 0.312$\pm$0.061
SVM(388) 0.034$\pm$0.01 0.103$\pm$0.017 0.034$\pm$0.014 0.392$\pm$0.071 0.077$\pm$0.012
Ackley(150, 150) 0.017$\pm$0.008 0.007$\pm$0.001 0.007$\pm$0.001 0.034$\pm$0.007 0.007$\pm$0.001
Ackley150(300, 150) 0.023$\pm$0.006 0.012$\pm$0.001 0.011$\pm$0.001 0.080$\pm$0.024 0.021$\pm$0.003
Hartmann6 (300, 6) 0.017$\pm$0.013 0.419$\pm$0.048 0.024$\pm$0.023 0.583$\pm$0.094 0.416$\pm$0.084
Rosenbrock(100, 100) 0.084$\pm$0.013 0.118$\pm$0.023 0.124$\pm$0.025 0.378$\pm$0.069 0.097$\pm$0.015
Rosenbrock(300, 100) 0.090$\pm$0.064 0.160$\pm$0.09 0.131$\pm$0.075 0.533$\pm$0.214 0.138$\pm$0.075
Stybtang(200, 200) 0.043$\pm$0.018 0.036$\pm$0.014 0.039$\pm$0.012 0.253$\pm$0.040 0.038$\pm$0.009
Table 4: Mean Squared Error(MSE) of each surrogate model. The results were averaged from five runs.
Figure 4: Optimization performance when using ARD Matérn trained with MAP and MLE, and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
4.5 Ablation Study and Failure Modes with Other Settings
Finally, we conducted a careful ablation study on the alternative settings to carry out standard BO. We found, however, those settings, such as using SE kernel and EI acquisition, often lead to inferior performance in high-dimensional optimization, which align with the a few failure cases reported in the literature [15, 20]. The detailed results and analysis are given in Section B of the Appendix. We summarize the favorable and unfavorable settings for high-dimensional optimization in Table 1.
@@ 5 Conclusion
We have presented a comprehensive investigation on standard BO (SBO) in high-dimensional optimization. We discovered that simply using Matérn kernel and UCB, SBO has already achieved the state-of-the-art performance. There seems no need for imposing additional structure assumptions for function learning and/or placing strong priors over the model parameters. In the future, we will continue applying SBO in more tasks, and seek for theoretical justifications and guarantees.
Appendix A More Experiment Details
@@ A.1 Definition of Synthetic Functions
Ackley. We used the Ackley function defined in [30],
@@
where each $x_{i}\in[0,1]$, ${\bf A}=[A_{ij}]$ and ${\bf P}=[P_{ij}]$ are two pre-defined matrices as in [30]. The global optimum is at ${\bf x}^{\dagger}=[0.20169,0.150011,0.476874,0.275332,0.6573]$.
@@
Figure 5: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in synthetic benchmarks. The results were averaged from 10 runs.
@@
Figure 6: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in real-world benchmarks. The results were averaged from 10 runs.
Figure 7: Ultimate optimization performance: Maximum Function Value obtained at the end of optimization loop, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel. The results were collect from 10 runs.
Figure 8: Runtime performance: Maximum Function Value obtained vs. Number of Steps in real-world benchmarks, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel. The results were averaged from 10 runs. The shaded region depicts the standard deviation.
A.2 Implementation
  • ALEBO. We used ALEBO implementation shared by the Adaptive Experimentation (AX) Platform (version 0.2.2). The source code is at https://github.com/facebook/Ax/blob/main/ax/models/torch/alebo.py.
  • HESBO. We used HESBO implementation of the original authors (https://github.com/aminnayebi/HesBO).
  • TURBO. We used the original TURBO implementation (https://github.com/uber-research/TuRBO)
  • SaasBO-NUTS. We used the original implementation of the authors (https://github.com/martinjankowiak/saasbo).
  • SaasBO-MAP. The SaasBO implementation available to the public does not include the version using MAP estimation. We therefore implemented this method based on the original paper [6]. All the hyperparameter settings follow exactly the same as the original paper.
  • RDUCB and Tree UCB. The implementation of both methods is publicly available at https://github.com/huawei-noah/HEBO/tree/master/RDUCB.
[…]
A.3 Computing Environment
@@
We conducted the experimental investigations on a large computer cluster equipped with Intel Cascade Lake Platinum 8268 chips.
Task ARD Matérn Matérn SaasBO Proj-GP Add-GP
Mopta08(124) 0.570$\pm$0.08 -0.277$\pm$0.101 0.560$\pm$0.064 -0.962$\pm$0.183 -0.089$\pm$0.105
Rover(60) -0.975$\pm$0.216 -0.647$\pm$0.158 -0.638$\pm$0.14 -0.995$\pm$0.116 -0.703$\pm$0.125
DNA(180) -0.256$\pm$0.261 -0.666$\pm$0.102 -0.183$\pm$0.158 -1.137$\pm$0.143 -0.679$\pm$0.143
SVM(388) 0.581$\pm$0.119 -0.179$\pm$0.042 0.650$\pm$0.12 -0.837$\pm$0.095 0.062$\pm$0.129
Ackley(150, 150) 0.428$\pm$0.222 0.633$\pm$0.067 0.574$\pm$0.068 -0.006$\pm$0.268 0.569$\pm$0.066
Ackley(300, 150) 0.232$\pm$0.155 0.252$\pm$0.057 0.279$\pm$0.077 -0.372$\pm$0.185 0.220$\pm$0.074
Hartmann6(300, 6) 1.200$\pm$0.169 -0.625$\pm$0.079 1.528$\pm$0.564 -0.927$\pm$0.084 -0.617$\pm$0.118
Rosenbrock(100, 100) 0.640$\pm$0.081 0.445$\pm$0.062 0.526$\pm$0.08 -0.466$\pm$0.182 0.521$\pm$0.057
Rosenbrock(300, 100) 0.518$\pm$0.225 0.121$\pm$0.155 0.357$\pm$0.21 -0.706$\pm$0.209 0.189$\pm$0.2
Stybtang(200, 200) 0.574$\pm$0.164 0.563$\pm$0.073 0.543$\pm$0.067 -0.458$\pm$0.069 0.582$\pm$0.072
@@
Table 5: Test log likelihood of each surrogate model. The results were averaged over five runs. Bolded are the best and second best results.
Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
ARD Matérn-MAP 7.39 9.99 7.81 9.77 13.55
ARD Matérn-MLE 11.23 9.84 5.76 14.34 32.16
ARD Matérn-NUTS 492.62 404.17 393.62 575.03 768.29
SaasBO-MAP 46.53 41.36 42.35 51.33 67.37
SaasBO-NUTS 985.76 375.06 229.07 1739.56 2269.50
@@
Figure 9: Optimization performance using MLE training, MAP training with a diffuse prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
Appendix B Ablation Study and Failure Modes With Other Settings
In this section, we conducted a careful ablation study to investigate alternative popular settings in executing standard BO. From these settings, we identified a few failure modes during high-dimensional optimization. We provided analysis and discussions accordingly.
B.1 GP Training
First, we examined MLE training with LBFGS. We found that, however, LBFGS often incurs numerical issues, and cannot reliably finish model training. This is consistent with the report in other works [12]. The issue might be because that without constraints (or regularization) over the parameters (such as the length-scales and amplitude), the exploration of LBFGS tends to be overly aggressive, which easily cause numerical problems. Hence, using stochastic optimization, like ADAM, might be a wise choice for MLE training.
Second, we examined MAP training with alternative prior choices over the length-scale parameters. In the main paper, we use a diffuse prior, Uniform(0.001, 30). In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular BO library BoTorch. To see how it is concentrated, in Fig. 10 we compare its probability density curve with other priors, including the aforementioned uniform prior and the diffuse Gamma priors over the amplitude and noise variance.
We found that employing a concentrated prior over the length-scale parameters will severely deteriorate the high-dimensional optimization performance. As shown in Fig 12 and 13, after placing the prior Gamma(3,6) over each length-scale parameter, both the ultimate and runtime optimization performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training. To figure out the reason, we further investigated the surrogate learning performance. Following the same setting as in Fig. 3, we examined the prediction alignment after applying the prior Gamma(3,6) over the length-scale parameters. As depicted in Fig. 11 (third and fourth column), this concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function. This is consistent with the report in the SaasBO paper [6] (see Fig.1 of that paper). The quantitative results shown in Table 8 and 9 further confirm the substantially worse prediction accuracy as compared with GP training with a diffuse prior. In addition, we notice that the previous work[21] have also tested standard BO using Gamma(3,6) as the prior over the length-scale parameters. In their high-dimensional optimization experiments (see Fig. 5 of that paper), standard BO performs badly. We used our setting (SBO-1 and SBO-2) to re-run one experiment that maximizes a high-dimensional polynomial function. It turns out that both SBO-1 and SBO-2 can surpass the best method I-BNN reported by that paper.
Together these results have shown that a concentrated prior can severely limit GP’s capacity for high-dimensional function estimation, leading to inferior performance in high-dimensional optimization.
Figure 10: Probability density curves. Gamma(2, 0.15) and Gamma(1.1, 0.5) are the priors for the amplitude and noise variance, respectively.
Figure 11: Alignment between the prediction (y-axis) and ground-truth function values (x-axis).
Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
LBFGS ADAM
Mopta08(124) 0.034$\pm$0.006 0.035$\pm$0.005 0.837$\pm$0.112 0.856$\pm$0.114
Rover(60) 0.327$\pm$0.078 0.323$\pm$0.075 0.78$\pm$0.154 0.819$\pm$0.115
DNA(180) 0.149$\pm$0.038 0.166$\pm$0.049 1.014$\pm$0.097 1.026$\pm$0.101
SVM(388) 0.034$\pm$0.01 0.039$\pm$0.01 1.057$\pm$0.066 1.06$\pm$0.067
Ackley(150, 150) 0.017$\pm$0.008 0.026$\pm$0.007 0.019$\pm$0.004 0.019$\pm$0.005
Ackley(300, 150) 0.023$\pm$0.006 0.07$\pm$0.047 0.431$\pm$0.09 0.459$\pm$0.092
Hartmann6(300, 6) 0.017$\pm$0.013 0.178$\pm$0.325 0.881$\pm$0.045 0.906$\pm$0.034
Rosenbrock(100, 100) 0.084$\pm$0.013 0.081$\pm$0.012 0.9$\pm$0.156 0.925$\pm$0.172
Rosenbrock(300, 100) 0.09$\pm$0.064 0.116$\pm$0.078 0.971$\pm$0.222 1.001$\pm$0.217
Stybtang(200, 200) 0.043$\pm$0.018 0.054$\pm$0.024 0.698$\pm$0.05 0.732$\pm$0.034
Table 8: Mean Squared Error(MSE) averaged from five runs. The best results are bolded.
Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
LBFGS ADAM
Mopta08(124) 0.57$\pm$0.08 0.542$\pm$0.104 -1.337$\pm$0.058 -1.715$\pm$0.212
Rover(60) -0.975$\pm$0.216 -1.062$\pm$0.286 -$1$1.322$\pm$0.073
DNA(180) -0.256$\pm$0.261 -0.388$\pm$0.35 -1.426$\pm$0.054 -2.026$\pm$0.183
SVM(388) 0.581$\pm$0.119 0.504$\pm$0.128 -1.453$\pm$0.038 -2.095$\pm$0.129
Ackley(150, 150) 0.428$\pm$0.222 0.14$\pm$0.218 0.451$\pm$0.088 0.445$\pm$0.087
Ackley(300, 150) 0.232$\pm$0.155 -0.257$\pm$0.528 -1.006$\pm$0.091 -1.037$\pm$0.087
Hartmann6(300, 6) 1.2$\pm$0.169 0.67$\pm$0.867 -1.344$\pm$0.029 -1.618$\pm$0.128
Rosenbrock(100, 100) 0.64$\pm$0.081 0.606$\pm$0.1 -1.308$\pm$0.083 -1.598$\pm$0.42
Rosenbrock(300, 100) 0.518$\pm$0.225 0.347$\pm$0.325 -1.411$\pm$0.134 -1.967$\pm$0.413
Stybtang(200, 200) 0.574$\pm$0.164 0.421$\pm$0.228 -1.251$\pm$0.017 -1.41$\pm$0.067
Table 9: Test log likelihood averaged from five runs. The best results are bolded.
Figure 12: Ultimate optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters. SBO-1 and SBO-2 use Uniform(0.001, 30).
Figure 13: Runtime optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters. SBO-1 and SBO-2 use Uniform(0.001, 30).
B.2 Alternative Acquisition Functions
Next, we examined two other popular acquisition functions used in BO: Expected Improvement (EI) and Thompson sampling (TS). At each step, EI computes the expected improvement upon the best sample so far,
$\text{EI}({\bf x})=(\mu({\bf x})-f({\bf x}^{*}))\Psi\left(\frac{\mu({\bf x})-f({\bf x}^{*})}{\sqrt{v({\bf x})}}\right)+\sqrt{v({\bf x})}\Phi\left(\frac{\mu({\bf x})-f({\bf x}^{*})}{\sqrt{v({\bf x})}}\right),$ (6)
where $\Psi(\cdot)$ and $\Phi(\cdot)$ are the CDF and PDF of the standard Gaussian distribution, ${\bf x}^{*}$ is the best sample that gives the largest function value among all the queries up to the current step, $\mu({\bf x})$ and $\sqrt{v({\bf x})}$ are the posterior mean and standard deviation of $f({\bf x})$ given current training data. We used BOTorch to maximize EI through L-BFGS. To conduct TS at each step, we used Sobol Sequence [29] to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior. We then selected the input with the largest function value as the next query. We denote this method by TS-3K. To explore more input candidates, we also employed the Lancos and conjugate gradient descent to approximate the posterior covariance matrix, and sampled 10K input candidates from Sobol Sequence. We denote this method as TS-10K. We used BOTorch implementation of both methods. We used ARD Matérn kernel and MLE training at each step.
The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig. 14 and 15, respectively. As we can see, in most cases, such as Mopta08 (128), SVM (388), DNA (180) and Rosenbrock (100, 100), both EI and TS perform worse than UCB, and also worse than the best baseline. On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst.
The limitation of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1]. A remedy is to use log-EI as the acquisition function. The failure of the TS might be due to the high-dimension of the input space. A few thousand candidates might not be sufficient to cover the whole space; as a result, the selected queries are inferior in indicating the best direction to optimize the target function.
Figure 14: Ultimate performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
Figure 15: Runtime performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
B.3 Using Square Exponential (SE) kernels
Third, we considered another widely used kernel in BO, namely, the SE kernel. We tested with ARD SE and SE kernel using a single length-scale parameter. We used MLE training and UCB $(\lambda=1.5)$ to run BO. As depicted in Fig. 17 and 18, with smaller optimization dimensions, such as Mopta08 (124), NAS201 (3) and DNA (180), the performance of SE is among the best. However, with higher dimensions, e.g., above 200, the performance of SE drops substantially, such as in SVM (388) and Ackley(300, 150).
We argue that the inferior performance might be due to that the vanilla SE kernel is not powerful enough to enable GP to well capture high-dimensional functions. Consider a single length-scale parameter. we observe that the value of the SE kernel is determined by the squared distance between the inputs,
$\kappa_{\text{SE}}({\bf x},{\bf x}^{\prime})=a\exp\left(-\eta\|{\bf x}-{\bf x}^{\prime}\|^{2}\right).$ (7)
With the growth of the input dimension $d$, the kernel function value decays much faster than the increase of the distance: a squared decay coupling with the exponential decay. The learning can be easily driven to adapt to a sparse, diagonal-dominant covariance matrix, which roughly treats each data point as independent. This can make the learning fail to capture the correlation between function values, and result in inferior learning performance. By contrast, the value of the Matérn kernel (see (1)) is dominated by the distance rather than the squared distance. Therefore, with the growth of the distance, the kernel function value decays much slower than the SE kernel. It is easier for the optimization to adjust the length-scale parameter so as to avoid “diagonal-like” covariance matrices. Consequently, the correlation between function values can be better captured in the high dimensional input space.
To verify our argument, we examined the GP learning performance with ARD SE kernel and ARD Matérn kernel. We used Rosenbrock(d,d) and Hartmann6(d, 6) as test cases. The former is a “full-rank” function, namely, all the $d$ input variables are relevant to the function value. The latter has a “low-rank” structure in that only the first 6 variables out of the $d$ input variables determines the function value. For each test case, we increased the input dimension from 50 to 300. We randomly generated 500 training examples and 100 test examples. We then trained each method with MLE, and examined the mean square error (MSE) and the test log likelihood. We repeated the training and test procedure for five times, and report the average MSE, average test log likelihood, and the standard deviations in Fig. 16.
As we can see, in both test cases, when the input dimension $d$ does not exceed 150, the performance of using SE and Matérn-5/2 is close, in terms of both MSE and log likelihood. The two curves almost overlap. However, for $d>150$ in both test scenarios, the MSE significantly escalates when employing the SE kernel, accompanied by a substantial decrease in the test log likelihood. These observations suggest a failure of the SE kernel in effectively learning the target function. In comparison, employing the Matérn kernel maintains a consistently favorable prediction performance even when $d>150$. On the Rosenbrock dataset, the MSE gradually increases while the test log likelihood decreases, following a roughly linear trend. This outcome is expected as, with the rise in input dimension $d$, more training examples are necessary to sustain comparable performance. Furthermore, on the Hartmann6 dataset, little variation is observed in both MSE and test log likelihood with increasing the input dimension when using the Matérn kernel. This suggests that the GP has effectively discerned the pertinent input variables and mitigated the influence of irrelevant ones. In summary, these results collectively underscore the advantage of employing Matérn kernels for learning high-dimensional functions. It is worth noting that the most recent work [12] designs a log-normal prior over each length-scale parameter to overcome the learning issue with the ARD SE kernel.
Figure 16: Learning high-dimension functions with different kernels. We trained on 500 points, and tested on 100 points. Results shown with mean and standard deviation over 5 runs.
Figure 17: Ultimate optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
Figure 18: Runtime optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
B.4 Different Choice of $\lambda$ in UCB
Finally, we examined how the exploration parameter $\lambda$ in UCB (see (4)) influences the performance in high-dimensional optimization. To this end, we used ARD Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}. We conducted MLE training via ADAM optimization. We show the ultimate and runtime optimization performance in Fig. 19 and 20, respectively. It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close. However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$. This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
Figure 19: Ultimate optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.
Figure 20: Runtime optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.